[Hallc_running] [New Logentry] Status Update

pooser at jlab.org pooser at jlab.org
Mon Feb 26 10:35:01 EST 2018

Logentry Text:
Beam during swing yesterday and owl this morning was quite ratty with ~20 trips / hr at times (https://logbooks.jlab.org/entry/3538786).  There was a ~2.5 hr down (https://logbooks.jlab.org/entry/3538641) from 2:15 - 4:45 am as a result of the trim rack W501B13 tripping off.  Once the issue was resolved beam was quickly restored to the Hall and production data taking continued.

At the beginning of day shift we successfully completed the kinematic point for the SHMS of 39 degrees at -2.5 GeV by successfully collecting production data for the LH2, LD2, and 10 cm Al dummy targets.  At ~7:45 am the SHMS momentum was reduced to -2.0 GeV (https://logbooks.jlab.org/entry/3538587) in order to facilitate the next kinematic point of interest for F2.  We plan to stay parked here for the next couple of days.

MCC began conducting beam studies today at 9 am, and will continue their studies until the end of day shift during which we will take parasitic beam for the first 4 hours.  When MCC begins their optics studies at 1 pm, we will take a controlled access to conduct various work in the hall.  We have been in contact with RadCon who reports that they will likely get to us some time around 1:30 pm. MCC will lock us up at 3:30 pm which results in mere 2 hour time window for any work in the hall to be conducted.

As a result of the reduced controlled access time window, Andy’s Y strainer installation on the SHMS power supply LCW lines has been rescheduled to the next controlled access.  When exactly the next controlled access will occur is unknown at the moment.  I will provide updates as soon as more information becomes available.

Work currently planned during at ~1:30 pm:
Joe Beaufait: diagnose (and potentially fix) HMS door/shutter breaker (https://logbooks.jlab.org/entry/3535292)
Howard Fenker: visually inspect HMS hut for causes for chamber ringing and 60 Hz noise
Steve Lassiter et. al.: Inspect HMS Q2 power supply via. visual and olfactive methods (thanks Simona :-p)

Special notices in effect for shift crews:
Trouble shooting HMS Q1 valves JT6 and JT8: https://logbooks.jlab.org/entry/3538239, https://logbooks.jlab.org/entry/3538243, https://logbooks.jlab.org/entry/3538254
Trouble shooting HMS Q2 power supply: https://logbooks.jlab.org/entry/3537306
Don't ramp HMS Q2 higher than 610A: https://logbooks.jlab.org/entry/3535714
TO Instructions for LH2 <--> LD2 switches : https://logbooks.jlab.org/entry/3534487

We will meet today in 200C at 1:30 pm (second floor of the counting house)
Link: https://bluejeans.com/543903672
Connecting directly from a room system?
1) Dial: or bjn.vc
2) Enter Meeting ID: 543 903 672
Just want to dial in on your phone?
1) Direct-dial with my iPhone or
+1.408.740.7256 (United States)
+1.888.240.2560 (US Toll Free)
+1.408.317.9253 (Alternate number)
(Global Numbers)
2) Enter Meeting ID: 543 903 672
3) Press #


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3538801
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180226/e84eca27/attachment-0001.html>

More information about the Hallc_running mailing list